OpenStudy (anonymous):

Solve 2x + 9 = 33 A. 7.5 B. 12 C. 21 D. 84

OpenStudy (anonymous):

a) subtract 9 from both sides b) divide both sides by 2

OpenStudy (anonymous):

2x + 9 = 33 - 9 = -9 2x=24 /2= /2 x=12

OpenStudy (ghostgate):

satellite73, is correct. First of all, subtract both sides of the equation by 9, what does that leave you with? 2x + 9 - 9 = 33 - 9? - What does that give you?

OpenStudy (anonymous):

The key here is to isolate the x by doing the same thing to both sides of the equation.

OpenStudy (ghostgate):

In the end you ended up doing all the work for him...

OpenStudy (♪chibiterasu):

Please do not give out direct answers, it is strictly against our Code of Conduct: http://openstudy.com/code-of-conduct